# Fundamentals and Advances in Transverse Thermoelectrics

### v1

@@ III.3 Performance of transverse thermoelectric refrigerator
In this subsection, we discuss the performance of a transverse thermoelectric refrigerator. For simplicity, we focus on the maximum attainable temperature difference of the device as a measure of its performance. In the previous two sections, we discarded the Joule heating effect as we used an approximation that is valid up to linear order in the temperature bias. By contrast, in order to investigate the maximum temperature difference, we need to take the Joule heating into account, and determine the temperature distribution in the refrigerator.
@@
where $c_{1}$ and $c_{2}$ are coefficients to be determined. Then, substituting Eq. (73) into Eq. (72) and setting $x=0$, we obtain
$ac_{1}+c_{2}+b=0.$
Besides, using the boundary condition $T(x=l_{x})=T_{\rm h}$, we have
$c_{1}l_{x}+\frac{c_{2}}{2}l_{x}^{2}=\Delta T,$
where $\Delta T=T_{\rm h}-T_{\rm c}$ as before. From these two equations, by discarding terms higher in $a^{2},ab$, the two coefficients $c_{1}$ and $c_{2}$ are approximately obtained as
$c_{1}$ $=$ $\frac{\Delta T}{l_{x}}+\frac{a}{2}\Delta T+\frac{b}{2}l_{x},$
$c_{2}$ $=$ $-a\frac{\Delta T}{l_{x}}-b,$
and the temperature distribution is given by
@@
Given the temperature distribution above, we are now in a position to calculate the rate of the heat removal from the cold reservoir ${\cal Q}_{\rm c}=Aq_{x}(x=0)$, where $A=l_{y}l_{z}$. Substituting Eq. (74) into Eq. (68) and multiplying $A$, we obtain
${\cal Q}_{\rm c}=-K\Delta T-\frac{R}{2}J_{y}^{2}-T_{\rm h}S_{yx}(l_{y}/l_{x})J_{y},$
where $J_{y}=l_{x}l_{z}\langle j_{y}\rangle_{\rm s}$, and $K$ and $R$ are defined below Eq. (47). We complete the square with respect to $J_{y}$ in the above equation as
${\cal Q}_{\rm c}=-K\Delta T-\frac{R}{2}\left(J_{y}-J_{y}^{*}\right)^{2}+\frac{T_{\rm h}^{2}S^{2}_{yx}}{2R}(l_{y}/l_{x})^{2},$
where $J_{y}^{*}=-{T_{\rm h}S_{yx}(l_{y}/l_{x})}/{R}$. Then, we see that the maximum value
${\cal Q}_{\rm c}=\frac{T_{\rm h}^{2}S^{2}_{yx}(l_{y}/l_{x})^{2}}{2R}-K\Delta T$
is obtained for $J_{y}=J_{y}^{*}$. Finally, the maximum temperature difference $\Delta T_{\rm max}$ is obtained for zero heat load ${\cal Q}_{\rm c}$ at the cold reservoir, yielding
@@
Since the temperature distribution is obtained in the same way, we now proceed exactly in the same manner as in the preceding subsection, and obtain the rate of the heat removal from the cold reservoir as
${\cal Q}_{\rm c}=-{K}^{\rm(a)}\Delta T-\frac{R^{\rm(b)}}{2}J_{y}^{2}-T_{\rm h}{S}^{\rm(a)}_{yx}(l_{y}/l_{x})J_{y},$
where ${K}^{\rm(a)}={\kappa}^{\rm(a)}_{xx}(l_{y}l_{z}/l_{x})$, $R^{\rm(b)}=\rho^{\rm(b)}_{yy}l_{y}/(l_{x}l_{z})$, and $J_{y}=l_{x}l_{z}\langle j_{y}\rangle_{\rm s}$. This is transformed as
@@
where $J_{y}^{*}=-{T_{\rm h}{S}^{\rm(a)}_{yx}(l_{y}/l_{x})}/{R^{\rm(b)}}$. From this result we see that ${\cal Q}_{\rm c}$ is maximized for $J_{y}=J_{y}^{*}$, yielding the following maximum temperature difference
@@ III.4 Performance of transverse active cooler
where $a=2S_{yx}\langle j_{y}\rangle_{\rm s}/\kappa_{xx}$ and $b=\rho_{yy}\langle j_{y}\rangle_{\rm s}^{2}/\kappa_{xx}$. Note that this expression is also obtained by interchanging $T_{\rm h}$ and $T_{\rm c}$ in Eq. (74). Then, substituting this temperature distribution into Eq. (68), we obtain the rate of heat removal from the hot reservoir ${\cal Q}_{\rm h}=l_{y}l_{z}q_{x}(x=0)$ as
${\cal Q}_{\rm h}=K\Delta T-\frac{R}{2}J_{y}^{2}-T_{\rm c}S_{yx}(l_{y}/l_{x})J_{y},$
where $J_{y}=l_{x}l_{z}\langle j_{y}\rangle_{\rm s}$, and $K$ and $R$ are defined below Eq. (47). Completing the square with respect to $J_{y}$ in the above equation, we obtain
${\cal Q}_{\rm h}=K\Delta T-\frac{R}{2}\big{(}J_{y}-J_{y}^{*}\big{)}^{2}+\frac{T_{\rm c}^{2}S_{yx}^{2}}{2R}(l_{y}/l_{x})^{2},$
where $J_{y}^{*}=-T_{\rm c}S_{yx}(l_{y}/l_{x})^{2}/2R$. Then, by setting $J_{y}=J_{y}^{*}$, the maximum heat removal from the hot reservoir ${\cal Q}_{\rm h}^{*}$ is obtained as
${\cal Q}_{\rm h}^{*}=K\Delta T+\frac{T_{\rm c}^{2}S_{yx}^{2}}{2R}(l_{y}/l_{x})^{2}.$
Now we introduce the heat current density $q_{\rm h}^{*}={\cal Q}^{*}_{\rm h}/(l_{y}l_{z})$, and rewrite the above equation as
$q_{\rm h}^{*}=\kappa_{xx}\left(\frac{\Delta T}{l_{x}}\right)+\frac{T_{\rm c}^{2}S_{yx}^{2}}{2\rho_{yy}l_{x}}.$
Then, defining ~$\widetilde{\nabla}_{x}T=\Delta T/l_{x}$ which is positive by definition, the above equation can be represented as
@@ IV.1 Transverse thermoelectric generator in adiabatic condition
$\widehat{\cal C^{\prime}}=\begin{pmatrix}\frac{-{S}^{\rm(a)}_{xy}T}{{K}^{\rm(a)}(l_{x}/l_{y})}&\frac{1}{{K}^{\rm(a)}}\\ -{R}^{\rm(a)}&\frac{+{S}^{\rm(a)}_{xy}}{{K}^{\rm(a)}(l_{x}/l_{y})}\end{pmatrix},$ (89)
where the determinant is also modified as
$\det\widehat{\cal C^{\prime}}=\frac{{R}^{\rm(a)}}{{K}^{\rm(a)}}(1-z_{xy}^{\rm(a)}T).$
Note that the adiabatic figure of merit $z_{xy}^{\rm(a)}$ in the above equation is defined by Eq. (41), but the adiabatic resistivity therein is redefined by Eq. (88). Consistent with this change, instead of Eq. (42), the adiabatic figure of merit for the present system is bounded, i.e.
$10\leq z^{\rm(a)}_{xy}T<$ 1 . (90)
@@ IV.2 Transverse thermoelectric generator in isothermal condition
$\widehat{\cal D^{\prime}}=\begin{pmatrix}S_{xy}T(l_{y}/l_{x})&K\\ -R&S_{xy}(l_{y}/l_{x})\end{pmatrix},$ (92)
where we used $S_{yx}=S_{xy}$. Then, the determinant also changes as
$\det\widehat{\cal D^{\prime}}=KR(1+z_{xy}^{\rm(i)}T),$
which means that, instead of Eq. (51), the isothermal figure of merit for the present system is unbounded, i.e.
@@ IV.3 Transverse thermoelectric refrigerator
then we perform replacement $j_{y}\to\langle j_{y}\rangle_{\rm s}$. By contrast, in the present case, we have the symmetry $S_{xy}=S_{yx}$ such that the coefficient $a$ vanishes, i.e.
$a=0$
for the off-diagonal Peltier device under the isothermal condition, while the other coefficient becomes
$b=\frac{\rho_{yy}j_{y}^{2}}{\kappa_{xx}}.$
Then, performing the same calculation as in Sec. III.3, the maximum attainable temperature difference is obtained as
@@
We next discuss the off-diagonal Peltier refrigerator under the adiabatic condition. Repeating the same calculation as in Sec. III.3, we find
$a=0,$
and the coefficient $b$ is given by
$b=\frac{\rho^{\rm(b)}_{yy}j_{y}^{2}}{\kappa^{\rm(a)}_{xx}},$
where $\rho^{\rm(b)}_{yy}$ is defined by Eq. (78). Then, the maximum temperature difference is calculated to be
@@ IV.4 Transverse active cooler
Finally, we discuss the performance of the active cooler based on the off-diagonal Peltier effect. Let us first consider the case of isothermal condition in the transverse direction. In this case, as in the previous section, we need to determine the temperature distribution by solving Eq. (72). Then, we obtain the temperature distribution expressed by Eq. (82), where $a=0$ and $b=\rho_{yy}j_{y}^{2}/\kappa_{xx}$ should be noted. Then, substituting this temperature distribution into Eq. (68), we obtain the rate of heat removal from the hot reservoir ${\cal Q}_{\rm h}=l_{y}l_{z}q_{x}(x=0)$ as
${\cal Q}_{\rm h}=K\Delta T-\frac{R}{2}J_{y}^{2}+T_{\rm h}S_{yx}(l_{y}/l_{x})J_{y},$
where $J_{y}=l_{x}l_{z}j_{y}$. Completing the square with respect to $J_{y}$ in the above equation, we obtain
${\cal Q}_{\rm h}=K\Delta T-\frac{R}{2}\big{(}J_{y}-J_{y}^{*}\big{)}^{2}+\frac{T_{\rm h}^{2}S_{yx}^{2}}{2R}(l_{y}/l_{x})^{2},$
where $J_{y}^{*}=-{T_{\rm h}S_{yx}}(l_{y}/l_{x})/2R$. Then, by setting $J_{y}=J_{y}^{*}$, the maximum heat removal from the hot reservoir ${\cal Q}_{\rm h}^{*}$ is obtained as
${\cal Q}_{\rm h}^{*}=K\Delta T+\frac{T_{\rm h}^{2}S_{yx}^{2}}{2R}(l_{y}/l_{x})^{2}.$
Now we define $q_{\rm h}^{*}={\cal Q}^{*}_{\rm h}/(l_{y}l_{z})$, and rewrite the above equation as
$q_{\rm h}^{*}=\kappa_{xx}\left(\frac{\Delta T}{l_{x}}\right)+\frac{T_{\rm h}^{2}S_{yx}^{2}}{2\rho_{yy}l_{x}}.$
Then, with the definition of ~$\widetilde{\nabla}_{x}T=\Delta T/l_{x}$, the above equation can be transformed to Eq. (83), in which the effective thermal conductivity for the present case is given by
@@ V Measurement of transverse thermopower
The transverse thermoelectric effects can be investigated by measuring an electric field or voltage in a sample in the transverse direction ($y$ direction) while applying a uniform temperature gradient in the longitudinal direction ($x$ direction). In the measurement of the transverse thermoelectric effects due to structural symmetry breaking, which is independent of a magnetic field $H$ and/or magnetization $M$, the positions of electrodes for measuring the voltage must be carefully set up because asymmetry in the electrode positions cause the longitudinal thermopower to be superimposed on the transverse thermopower. It is easy to separate the transverse thermoelectric effects due to time-reversal symmetry breaking, which show an odd dependence on $H$ and/or $M$ in the $z$ direction, from the longitudinal thermopower. Here, the pure transverse thermopower can be extracted by measuring the $H$ dependence of the transverse voltage $V$ and calculating $(V(+H)-V(-H))/2$ with $V(+H)$ and $V(-H)$ respectively being $V$ obtained when positive and negative magnetic fields are applied. In the case of the longitudinal thermoelectric effects, where the temperature gradient and electric field are parallel, the voltage per temperature difference and the electric field per temperature gradient are the same value. In contrast, in the case of the transverse thermoelectric effects, these are different values depending on the aspect ratio of the sample. Therefore, to discuss the performance of the transverse thermoelectric effects, the electric field per temperature gradient, i.e. the transverse thermopower, must be compared.
@@ V.1 Measurement configuration for Nernst effect in thin films
In the fields of spin caloritronics and topological materials science, ANE has been widely measured not only in bulk materials but also in thin film devices. In contrast to simple measurements for bulk materials, there are two experimental configurations for thin films because of their huge difference between the in-plane and out-of-plane dimensions, and their characteristics need to be understood [34, 54]. One configuration is the perpendicularly magnetized (PM) configuration, where ${\bf H}$ ($\nabla T$) is applied along the out-of-plane (in-plane) direction [Fig. 6(a)]. The PM configuration is widely used for the quantitative estimation of $S_{\rm ANE}$ since the magnitude of $\nabla T$ can be precisely measured by various experimental techniques including thermometers attached on hot and cold sides of a substrate or heat baths [54], on-chip thermometers grown on the substrate [39], and an infrared camera [41, 42, 44, 45, 55]. Thus, by estimating the ANE-induced electric field through the extraction of the zero-field intercept of the $H$ dependence of the transverse electric field $E_{y}$ and by normalizing it by $\nabla T$, one can determine $S_{\rm ANE}$. However, the PM configuration often requires the application of large $H$ to align the magnetization of films along the out-of-plane direction to overcome the strong demagnetization field. The other configuration is the in-plane magnetized (IM) configuration, where ${\bf H}$ ($\nabla T$) is applied along the in-plane (out-of-plane) direction [Fig. 6(b)]. The IM configuration is suitable for thermal energy harvesting and heat flux sensing, discussed later, as ANE-based thermoelectric generation works simply by forming films onto heat sources. However, the IM configuration is not suitable for estimating $S_{\rm ANE}$ quantitatively because the temperature difference between the top and bottom of thin films is hard to be quantified (note that it is possible to quantify $S_{\rm ANE}$ even in the IM configuration by measuring the out-of-plane thermal conductivity of thin films using the time-domain thermoreflectance method [145]). Comparing the transverse thermopower in the PM and IM configurations is also used to separate the contribution of ANE from that of spin-Seebeck-driven ISHE, which appears only in the IM configuration due to the symmetry of ISHE [146, 147]. Note that the situations discussed in this subsection are applicable to ONE in thin films.
Figures 6(c) and 6(d) show an example of the experimental results of ANE measured in the PM and IM configurations, respectively, for a 50-nm-thick Co2MnGa thin film, epitaxially grown on a single-crystalline MgO (001) substrate [34]. In both the configurations, the observed transverse thermoelectric signals showed an odd dependence on $H$ and its magnitude saturated in the high-$H$ region, where the magnetization of the Co2MnGa film saturated. This is a typical behavior of ANE. In the IM configuration, a large ANE-induced thermopower was observed to appear even at zero field because the strong demagnetization field in the out-of-plane direction stabilized the remanent magnetization in the in-plane direction [Fig. 6(d)], a situation preferable for practical applications. In contrast, a tiny ANE-induced thermopower was observed at zero field in the PM configuration [Fig. 6(c)]. To generate the ANE-induced thermopower in the absence of an external magnetic field in the PM configuration, it is necessary to use a magnetic thin film having large perpendicular magnetic anisotropy [18, 52].

### v2

@@ III.3 Performance of transverse thermoelectric refrigerator
In this subsection, we consider the performance of a transverse thermoelectric refrigerator/heat pump. Here, as mentioned in Introduction, we focus only on the cooling mode and discuss the maximum attainable temperature difference of the device as a measure of its performance. Note that the thermoelectric refrigerator discussed here is different from compressor-based refrigerators and magnetocaloric refrigerators [139]. In the previous two sections, we discarded the Joule heating effect as we used an approximation that is valid up to linear order in the temperature bias. By contrast, in order to investigate the maximum temperature difference, we need to take the Joule heating into account, and determine the temperature distribution in the refrigerator.
@@
where $c_{1}$ and $c_{2}$ are coefficients to be determined. Then, substituting Eq. (73) into Eq. (72) and setting $x=0$, we obtain $ac_{1}+c_{2}+b=0$. Besides, using the boundary condition $T(x=l_{x})=T_{\rm h}$, we have $c_{1}l_{x}+\frac{c_{2}}{2}l_{x}^{2}=\Delta T$, where $\Delta T=T_{\rm h}-T_{\rm c}$ as before. From these two equations, by discarding terms higher in $a^{2},ab$, the two coefficients $c_{1}$ and $c_{2}$ are approximately obtained as $c_{1}={\Delta T}/{l_{x}}+({a}/{2})\Delta T+({b}/{2})l_{x}$ and $c_{2}=-a{\Delta T}/{l_{x}}-b$, and the temperature distribution is given by
@@
Given the temperature distribution above, we are now in a position to calculate the rate of the heat removal from the cold reservoir ${\cal Q}_{\rm c}=Aq_{x}(x=0)$, where $A=l_{y}l_{z}$. Substituting Eq. (74) into Eq. (68) and multiplying $A$, we obtain ${\cal Q}_{\rm c}=-K\Delta T-({R}/{2})J_{y}^{2}-T_{\rm h}S_{yx}(l_{y}/l_{x})J_{y}$, where $J_{y}=l_{x}l_{z}\langle j_{y}\rangle_{\rm s}$, and $K$ and $R$ are defined below Eq. (47). We complete the square with respect to $J_{y}$ in the above equation as ${\cal Q}_{\rm c}=-K\Delta T-({R}/{2})\left(J_{y}-J_{y}^{*}\right)^{2}+({T_{\rm h}^{2}S^{2}_{yx}}/{2R})(l_{y}/l_{x})^{2}$, where $J_{y}^{*}=-{T_{\rm h}S_{yx}(l_{y}/l_{x})}/{R}$. Then, we see that the maximum value ${\cal Q}_{\rm c}=({T_{\rm h}^{2}S^{2}_{yx}(l_{y}/l_{x})^{2}})/({2R})-K\Delta T$ is obtained for $J_{y}=J_{y}^{*}$. Finally, the maximum temperature difference $\Delta T_{\rm max}$ is obtained for zero heat load ${\cal Q}_{\rm c}$ at the cold reservoir, yielding
@@ III.4 Performance of transverse active cooler
where $a=2S_{yx}\langle j_{y}\rangle_{\rm s}/\kappa_{xx}$ and $b=\rho_{yy}\langle j_{y}\rangle_{\rm s}^{2}/\kappa_{xx}$. Note that this expression is also obtained by interchanging $T_{\rm h}$ and $T_{\rm c}$ in Eq. (74). Then, substituting this temperature distribution into Eq. (68), we obtain the rate of heat removal from the hot reservoir ${\cal Q}_{\rm h}=l_{y}l_{z}q_{x}(x=0)$ as ${\cal Q}_{\rm h}=K\Delta T-({R}/{2})J_{y}^{2}-T_{\rm c}S_{yx}(l_{y}/l_{x})J_{y}$, where $J_{y}=l_{x}l_{z}\langle j_{y}\rangle_{\rm s}$, and $K$ and $R$ are defined below Eq. (47). Completing the square with respect to $J_{y}$ in the above equation, we obtain ${\cal Q}_{\rm h}=K\Delta T-({R}/{2})\big{(}J_{y}-J_{y}^{*}\big{)}^{2}+({T_{\rm c}^{2}S_{yx}^{2}}/{2R})(l_{y}/l_{x})^{2}$, where $J_{y}^{*}=-T_{\rm c}S_{yx}(l_{y}/l_{x})^{2}/2R$. Then, by setting $J_{y}=J_{y}^{*}$, the maximum heat removal from the hot reservoir ${\cal Q}_{\rm h}^{*}$ is obtained as ${\cal Q}_{\rm h}^{*}=K\Delta T+({T_{\rm c}^{2}S_{yx}^{2}}/{2R})(l_{y}/l_{x})^{2}$. Now we introduce the heat current density $q_{\rm h}^{*}={\cal Q}^{*}_{\rm h}/(l_{y}l_{z})$, and rewrite the above equation as $q_{\rm h}^{*}=\kappa_{xx}\left(({\Delta T}/{l_{x}})\right)+({T_{\rm c}^{2}S_{yx}^{2}})/({2\rho_{yy}l_{x}})$. Then, defining ~$\widetilde{\nabla}_{x}T=\Delta T/l_{x}$ which is positive by definition, the above equation can be represented as
@@ IV.1 Transverse thermoelectric generator in adiabatic condition
$\widehat{\cal C^{\prime}}=\begin{pmatrix}\frac{-{S}^{\rm(a)}_{yx}T}{{K}^{\rm(a)}(l_{x}/l_{y})}&\frac{1}{{K}^{\rm(a)}}\\ -{R}^{\rm(a)}&\frac{+{S}^{\rm(a)}_{yx}}{{K}^{\rm(a)}(l_{x}/l_{y})}\end{pmatrix},$ (89)
where the determinant is also modified as $\det\widehat{\cal C^{\prime}}=({{R}^{\rm(a)}}/{{K}^{\rm(a)}})(1-z_{xy}^{\rm(a)}T)$. Note that the adiabatic figure of merit $z_{xy}^{\rm(a)}$ in the above equation is defined by Eq. (41), but the adiabatic resistivity therein is redefined by Eq. (88). Consistent with this change, instead of Eq. (42), the adiabatic figure of merit for the present system is bounded, i.e.
$0\leq z^{\rm(a)}_{xy}T<1$ (90)
@@ IV.2 Transverse thermoelectric generator in isothermal condition
$\widehat{\cal D^{\prime}}=\begin{pmatrix}S_{yx}T(l_{y}/l_{x})&K\\ -R&S_{yx}(l_{y}/l_{x})\end{pmatrix},$ (92)
where we used $S_{yx}=S_{xy}$. Then, the determinant also changes as $\det\widehat{\cal D^{\prime}}=KR(1+z_{xy}^{\rm(i)}T)$, which means that, instead of Eq. (51), the isothermal figure of merit for the present system is unbounded, i.e.
@@ IV.3 Transverse thermoelectric refrigerator
then we perform replacement $j_{y}\to\langle j_{y}\rangle_{\rm s}$. By contrast, in the present case, we have the symmetry $S_{xy}=S_{yx}$ such that the coefficient $a$ vanishes, i.e. $a=0$ for the off-diagonal Peltier device under the isothermal condition, while the other coefficient becomes $b={\rho_{yy}j_{y}^{2}}/{\kappa_{xx}}$. Then, performing the same calculation as in Sec. III.3, the maximum attainable temperature difference is obtained as
@@
We next discuss the off-diagonal Peltier refrigerator under the adiabatic condition. Repeating the same calculation as in Sec. III.3, we find $a=0$ and $b={\rho^{\rm(b)}_{yy}j_{y}^{2}}/{\kappa^{\rm(a)}_{xx}}$, where $\rho^{\rm(b)}_{yy}$ is defined by Eq. (78). Then, the maximum temperature difference is calculated to be
@@ IV.4 Transverse active cooler
Finally, we discuss the performance of the active cooler based on the off-diagonal Peltier effect. Let us first consider the case of isothermal condition in the transverse direction. In this case, as in the previous section, we need to determine the temperature distribution by solving Eq. (72). Then, we obtain the temperature distribution expressed by Eq. (82), where $a=0$ and $b=\rho_{yy}j_{y}^{2}/\kappa_{xx}$ should be noted. Then, substituting this temperature distribution into Eq. (68), we obtain the rate of heat removal from the hot reservoir ${\cal Q}_{\rm h}=l_{y}l_{z}q_{x}(x=0)$ as ${\cal Q}_{\rm h}=K\Delta T-({R}/{2})J_{y}^{2}+T_{\rm h}S_{yx}(l_{y}/l_{x})J_{y}$, where $J_{y}=l_{x}l_{z}j_{y}$. Completing the square with respect to $J_{y}$ in the above equation, we obtain ${\cal Q}_{\rm h}=K\Delta T-({R}/{2})\big{(}J_{y}-J_{y}^{*}\big{)}^{2}+({T_{\rm h}^{2}S_{yx}^{2}}/{2R})(l_{y}/l_{x})^{2}$, where $J_{y}^{*}=-{T_{\rm h}S_{yx}}(l_{y}/l_{x})/2R$. Then, by setting $J_{y}=J_{y}^{*}$, the maximum heat removal from the hot reservoir ${\cal Q}_{\rm h}^{*}$ is obtained as ${\cal Q}_{\rm h}^{*}=K\Delta T+({T_{\rm h}^{2}S_{yx}^{2}}/{2R})(l_{y}/l_{x})^{2}$. Now we define $q_{\rm h}^{*}={\cal Q}^{*}_{\rm h}/(l_{y}l_{z})$, and rewrite the above equation as $q_{\rm h}^{*}=\kappa_{xx}\left({\Delta T}/{l_{x}}\right)+({T_{\rm h}^{2}S_{yx}^{2}}/{2\rho_{yy}l_{x}})$. Then, with the definition of ~$\widetilde{\nabla}_{x}T=\Delta T/l_{x}$, the above equation can be transformed to Eq. (83), in which the effective thermal conductivity for the present case is given by
@@ V Measurement of transverse thermopower
[…]
@@ V.1 Measurement configuration for Nernst effect in thin films
In the fields of spin caloritronics and topological materials science, ANE has been widely measured not only in bulk materials but also in thin film devices. In contrast to simple measurements for bulk materials, there are two experimental configurations for thin films because of their huge difference between the in-plane and out-of-plane dimensions, and their characteristics need to be understood [37, 57]. One configuration is the perpendicularly magnetized (PM) configuration, where ${\bf H}$ ($\nabla T$) is applied along the out-of-plane (in-plane) direction [Fig. 6(a)]. The PM configuration is widely used for the quantitative estimation of $S_{\rm ANE}$ since the magnitude of $\nabla T$ can be precisely measured by various experimental techniques including thermometers attached on hot and cold sides of a substrate or heat baths [57], on-chip thermometers grown on the substrate [42], and an infrared camera [44, 45, 47, 48, 58]. Thus, by estimating the ANE-induced electric field through the extraction of the zero-field intercept of the $H$ dependence of the transverse electric field $E_{y}$ and by normalizing it by $\nabla T$, one can determine $S_{\rm ANE}$. However, the PM configuration often requires the application of large $H$ to align the magnetization of films along the out-of-plane direction to overcome the strong demagnetization field. The other configuration is the in-plane magnetized (IM) configuration, where ${\bf H}$ ($\nabla T$) is applied along the in-plane (out-of-plane) direction [Fig. 6(b)]. The IM configuration is suitable for thermal energy harvesting and heat flux sensing, discussed later, as ANE-based thermoelectric generation works simply by forming films onto heat sources. However, the IM configuration is not suitable for estimating $S_{\rm ANE}$ quantitatively because the temperature difference between the top and bottom of thin films is hard to be quantified (note that it is possible to quantify $S_{\rm ANE}$ even in the IM configuration by measuring the out-of-plane thermal conductivity of thin films using the time-domain thermoreflectance method [156]). Comparing the transverse thermopower in the PM and IM configurations is also used to separate the contribution of ANE from that of SSE-driven ISHE, which appears only in the IM configuration due to the symmetry of ISHE [157, 158]. Note that the situations discussed in this subsection are applicable to ONE in thin films.
[…]